OpenStudy (anonymous):

7 notebooks cost \$6.16 Which equation would help determine the cost of 4 notebook

OpenStudy (moazzam07):

3.54

OpenStudy (moazzam07):

sorry 3.52

OpenStudy (moazzam07):

$3.52 is the value of 4 notebooks

OpenStudy (kaitlyn_nicole):

find the cost of one by dividing the 6.16 and 7 then take that amount and times it by 4.

OpenStudy (moazzam07):

6.16 divided by 7 = 0.88 .88 times 4 = 3.52

OpenStudy (anonymous):

i got to write a proportion

OpenStudy (moazzam07):

@Black_ninja123 0.94? is wrong

OpenStudy (texaschic101):

or you could set it up as a proportion... 7/6.16 = 4/x --- 7 notebooks for 6.16 = 4 notebooks for $x

OpenStudy (black_ninja123):

Yea 0.94 is wrong sorry

OpenStudy (kaitlyn_nicole):

x=(7/6.16)x4

OpenStudy (texaschic101):

7/6.16 = 4/x --- cross multiply 7x = (6.16)(4) can you finish and solve for x

OpenStudy (anonymous):

uhhhg i dont understand so many replies

OpenStudy (anonymous):

its just writing proportions

OpenStudy (texaschic101):

proportions are nothing but equivalent fractions

OpenStudy (texaschic101):

7/6.16 = 4/x --- 7 notebooks for 6.16 = 4 notebooks for x dollars CROSS MULTIPLY (7)(x) = (6.16)(4) 7x = 24.64 x = 24.64/7 x = 3.52

OpenStudy (kaitlyn_nicole):

x=(4/6.80) x 33

OpenStudy (anonymous):

huh?

OpenStudy (anonymous):

4 glue sticks cost \$6.80 Which equation would help determine the cost of 3 glue sticks?

OpenStudy (kaitlyn_nicole):

x=(4/6.80) x 3 meaning that bc the 4/6.80 is in parenthesis then you do that first to find the cost of one, then you take the quotient of that and times it by three to find the cost.

OpenStudy (anonymous):

i dont understand? wht?

OpenStudy (moazzam07):

what grade are you in?

OpenStudy (anonymous):

hmmmmmmm lets see uhhh not ur business